# EvalLM: Interactive Evaluation of Large Language Model Prompts on User-Defined Criteria

### v1

@@ 6.2.1. Procedure
We sampled two evaluations by the Specific- Criteria condition for each of the 19 tasks, resulting in a total of 38 output pairs evaluated and 194 criteria-wise evaluations. Each evaluation was annotated for the presence of any errors regarding five criteria: (1) logical (i.e., the explanation presents logical and coherent arguments and justifications), (2) faithful (i.e., the explanation does not hallucinate content that does not exist in the outputs), (3) independent (i.e., the explanation does not assess other criteria or aspects not described in the evaluation criterion), (4) evidential (i.e., the evidence extracted is relevant to the explanation), and (5) score aligned (i.e., the final score aligns with the explanation provided). We recruited two annotators, who had previous experience grading written assignments, and instructed them to mark these errors even if only part of the explanation presented the error. Since there is a large class imbalance (i.e., most explanations have no errors), we considered an explanation to have errors if at least one of the evaluators marked an error.
@@ 6.2.2. Results
Overall, we found that the explanations were mostly free of issues: 91.4% of the explanations were logical, 99.1% were faithful, 84.2% were independent, 100% provided relevant evidence, and 98.6% were aligned with their scores. We qualitatively reviewed erroneous cases and observed that the LLM evaluations frequently failed to produce independent explanations as they would contrast outputs on their level of detail despite the criterion not assessing this. In terms of logic, the evaluations struggled to assess creativity and could be too superficial in their interpretations. For example, a story was considered novel for describing a character as “the last surviving member of a wizard family”, and news article headlines were considered to address ethical dilemmas since they explicitly included the phrase “ethical concerns”. For faithfulness, there were cases where the evaluations could not accurately measure the length of the outputs. Finally, for score alignment, there were cases where the explanations did not mention that one output was better than the other but provided one output with a higher score. These results show that GPT-4 is able to produce relatively sensible explanations for its evaluations, but may at times be limited by the model’s bias towards detail and its logical capabilities.
@@ 7. User Study
To understand how the EvalLM affects the prompt iteration process when compared to following the current practice of designers, we conducted a within-subjects study where we compared EvalLM to a baseline where participants manually evaluate outputs. In this study, we aimed to answer the following research questions:
RQ1. Can EvalLM aid designers in deciding on how to revise their prompts and verifying the effectiveness of these revisions?
@@ 7.1.1. Participants
We recruited 12 participants through posts on online forums within our institution. All of these participants reported having extensive experiences with prompting: nine had designed prompts for research-based applications, one designed prompts for toy projects, and two described frequent use of LLMs for productivity. Regarding the length of their experiences, four had between 1 to 3 months of experience, three had 3 to 6 months of experience, four had 6 to 12 months of experience, and two had 1 to 2 years of experience. Participants were compensated with approximately 60 USD (80,000 KRW) for the 2-hour study.
@@ 7.1.2. Conditions
During the study, participants designed prompts and evaluated their outputs for two given tasks. For each task, participants either used EvalLM in one of two conditions: Assist or Manual. The Assist condition was the full EvalLM interface, while the Manual condition was the EvalLM interface without the evaluation assistant or the criteria review tool. In the Manual condition, participants defined their own criteria or selected from the dictionary, and then evaluated each data sample by choosing which output won for these criteria. This condition supports designers’ common practices—according to our formative interviews—where they copy generated outputs into a spreadsheet to manually check which criteria were satisfied by each output.
@@ 7.1.3. Tasks
All of the participants designed prompts for the same two tasks. As our work focuses on supporting the design and evaluation of prompts for novel tasks, we adapted tasks from recently proposed LLM-powered HCI systems. Specifically, we chose the following two tasks: (1) write an example that can explain a piece of scientific information to a young child—based on Lee et al.’s DAPIE system (Lee et al., 2023)—and (2) ideate a list of alternative angles for a news story—based on Peridis et al.’s AngleKindling system (Petridis et al., 2023). We chose these two tasks as they target a specific user population (i.e., children and reporters) and no significant expertise is needed to understand the task outputs. Additionally, these two tasks are different in terms of their goal (i.e., explaining vs. brainstorming), and how the output relates to the input (i.e., transforming vs. expanding on the information).
@@ 7.1.4. Procedure
Participants were asked to sign the informed consent form prior to the study. After a brief introduction to the study, participants answered a pre-task survey. After a walkthrough of the first interface, participants used this interface to design a prompt for the first task.444The order of conditions and tasks were counterbalanced to mitigate ordering effects. Participants were asked to envision themselves as a developer at a startup that is building an application that performs the given generation task. Their goal was to design a prompt that performed better than an initial prompt that their team had designed and demonstrate its performance on diverse data samples. Participants could flexibly decide on the criteria that they would be improving and evaluating, but were asked to ensure that their final criteria set was: (1) exhaustively comprehensive (i.e., assess all factors that are important for the task), (2) mutually exclusive (i.e., minimal redundancies between criteria), and (3) clear (i.e., clearly described for others to understand what is assessed). After this explanation, participants performed the first task for 35 minutes. After the task, they responded to a post-task survey and, through a semi-structured interview, we asked them about how they defined their criteria, evaluated outputs, and revised their prompts during this task. Then, participants were provided with a walkthrough of the second interface and used this interface to perform the second task for 35 minutes. After the task, participants responded to the post-task survey and we concluded the study with a semi-structured interview about participants’ experiences in the second task, and the differences between their experiences in the first and second task.
@@ 7.1.5. Measures
For qualitative data, we transcribed the semi-structured interviews and coded them through a thematic analysis. For quantitative data, we analyzed participants’ responses to the two post-task surveys. These surveys asked participants to rate, on a seven-point Likert scale, their self-confidence in designing prompts and evaluating outputs, and their self-perceived experience interacting with the system (e.g., how successful and collaborative it was) based on questions from Wu et al. (Wu et al., 2022). We also asked participants to self-rate their final criteria on how exhaustively comprehensive, mutual exclusive, and clear they were. Finally, participants rated their self-perceived workload on five questions from the NASA-TLX questionnaire, excluding the “Physical Demand” question. We include the detailed survey questions in Appendix D. For these Likert scale ratings, we analyzed them through the non-parametric Wilcoxon signed-rank test. Additionally, we analyzed participants’ interaction logs to measure the number of (1) unique prompts tested, (2) criteria changes (e.g., edit, add, delete), and (3) unique outputs evaluated, where partial evaluations (i.e., only a subset of criteria were evaluated) were counted equally. We also analyzed how participants created and revised their final set of criteria. For all these measures, we conducted a Shapiro-Wilk test to determine if the data was parametric or non-parametric, and then used a paired t-test (if parametric) and a Wilcoxon signed-rank test (if non-parametric).
@@
7.2. Results
Our results demonstrated that EvalLM allowed participants to “debug” their prompts by quickly checking the effect of prompt changes on multiple criteria and samples. Beyond checking the effect of revisions, we observed that EvalLM provided “feedback” to participants, which helped them ideate on how to revise their prompts. Regarding criteria, EvalLM encouraged participants to revise their criteria more, which in turn allowed them to consider diverse values and perspectives around the given tasks. In this section, we describe findings on how participants evaluated outputs and revised their prompts in §7.2.1 and §7.2.1 (RQ1), how they defined their criteria in §7.2.1 (RQ2), their trust for the evaluation assistant in §7.2.1 (RQ3), and their overall perceived workload in §7.2.1. For each of these, we first describe relevant quantitative findings and then qualitative insights.
7.2.1. Evaluating Outputs
Overall, participants had higher self-confidence in their ability to evaluate prompt outputs with the Assist condition (Assist=6.71 ± 0.40, Manual=4.96 ± 0.72, z=9.23, p¡0.001). Participants revealed that this was due to how the Assist condition shouldered the burden of examining outputs, which enabled them to evaluate prompts more comprehensively. Specifically, with Assist, participants were able to evaluate a larger number of unique outputs (Assist=20.42 ± 14.46, Manual=10.08 ± 6.27, z=8.00, p=0.03). Several participants noted how this helped them look at the “bigger picture” (P1) and understand “how [a prompt] will work at a larger scale” (P5).
By facilitating evaluations, the Assist condition also allowed participants to evaluate more criteria and assess their prompts’ performance on diverse dimensions, without worrying about the cost involved. For example, P9 mentioned that they were encouraged to “think of more aspects that they wanted to evaluate” and P2 mentioned that they “just kept criteria […] to check them just in case.” In contrast, due to the effort involved, participants would frequently only evaluate samples on a subset of their criteria when using Manual—on average, 46.7% (SD=31.2%) of all evaluated samples were partially evaluated. Besides evaluating outputs for them, several participants also mentioned how the Assist condition made it easier for them to evaluate the outputs themselves. For example, both P4 and P6 mentioned how it could be challenging to “tell what is different by simply looking at [outputs]” (P5) and “to see why one output might be better or not” (P4). However, as the evaluation assistant explained its evaluations and highlighted sections of the outputs, it was “easier to tell outputs apart” (P6) and actually assess these outputs.
As the Assist condition supported faster evaluation of prompts on various samples and criteria, participants mentioned how they used the evaluation assistant as a “debugger” (P8). The evaluation assistant helped participants to check “what the prompt is already satisfying” (P8) and “where it’s lacking” (P9). This in turn allowed participants to decide on what aspects need the most improvement and focus their iterations on them. For example, P11 mentioned, “The [system] tells me that my prompt is worse on simplicity so that means that my prompt is not communicating this clearly and I should focus on fixing this.” Then, participants mentioned how they could check that these revisions had the desired effect by evaluating the outputs generated from the revised prompts.
@@ 7.2.2. Revising Prompts
Besides aiding participants in checking revisions, the Assist condition also helped participants in thinking about how to revise their prompts. Participants felt that the Assist condition helped them think about how to complete the task better (Assist=6.83 ± 0.39, Manual=5.67 ± 1.15, z=0.00, p=0.01) and that they were collaborating with the system to design the prompts (Assist=6.58 ± 0.51, Manual=4.58 ± 1.98, z=1.50, p¡0.01) (Fig. 8, left). Additionally, we found that participants with the Assist condition reached a similar level of satisfaction in their prompts (Assist=5.67 ± 1.61, Manual=5.17 ± 1.47, z=20.00, p=0.24) (“Match Goal” in Fig. 8) by testing a smaller number of prompt variations (Assist=5.00 ± 3.13, Manual=8.42 ± 4.76, z=1.50, p=0.01). While participants also reported a higher self-confidence on their ability to design and improve prompts with Assist, this difference was not significant (Assist=5.63 ± 1.03, Manual=5.00 ± 0.95, z=11.00, p=0.17).
In the Manual condition, participants mentioned how they felt like they were “the only brain thinking about [the task]” (P10). In contrast, as the Assist condition explains its evaluations, participants felt that the system was providing them with “feedback on what to improve” (P2) and could help them set the “direction” for their prompts (P7, P9). Several participants found the explanations to be particularly useful as they presented them with “diverse opinions” (P1) that they “couldn’t think about” (P6). For example, P5 purposefully increased the number of evaluation trials and checked the explanation for each trial to “see all the differences [in opinion] and then change my prompt to satisfy all of these.” In this sense, the Assist condition could allow participants to gain a more holistic understanding of how to improve their prompts. In turn, this understanding could help participants to make more effective prompt revisions—as reflected by the finding that participants tested less prompt changes in the Assist condition.
These explanations, however, could also lead to participants feeling less in control and more aware of their prompt’s weaknesses. Out of all participants, only P2 mentioned preferring the Manual condition. She described how she incorporated her “own diverse ideas” into the prompt in Manual, but, in Assist, she “kept worrying about the evaluations” as she focused on incorporating feedback but, at times, not seeing improvements in the evaluations. Even participants that preferred the Assist condition mentioned how the explanations could occasionally be limited as evaluations would frequently result in ties or provide high ratings to both outputs and they “wouldn’t know what to improve.” Since the feedback focused on the outputs rather than the prompts, several participants also noted how they would have preferred if the Assist condition would “automatically suggest ways to improve prompts.” As an alternative solution, P3 and P10 mentioned how they would add “keywords” from the explanations into their prompts.
7.2.3. Defining Criteria
Participants had a similar number of final criteria at the end of the tasks (Assist=4.25 ± 1.29, Manual=4.00 ± 1.28, t=10.50, p=0.55). In terms of their evaluation criteria, participants felt that their criteria in the Assist condition were more comprehensive (Assist=6.00 ± 1.48, Manual=4.75 ± 1.48, t=2.07, p=0.06), although with marginal significance, and clearer (Assist= 6.42 ± 0.67, Manual=4.92 ± 1.44, t=3.59, p¡0.01) than in the Manual condition (Fig. 8, right). Additionally, participants made more changes to their criteria with the Assist condition (Assist=22.67 ± 9.17, Manual=13.33 ± 8.64, t=2.29, p=0.04). For participants’ final criteria, there was no significant difference in the proportion of criteria created from scratch (Assist=30.1% ± 34.5%, Manual=16.67% ± 18.8%, w=15.0, p=0.37) or through the criteria dictionary (Assist= 77.78% ± 20.5%, Manual=69.93% ± 34.50%, w=22.0, p=0.95). However, in the Assist condition, a significantly higher proportion of the dictionary-based criteria had been edited (Assist=75.8% ± 24.1%, Manual=17.4% ± 34.5%, w=4.000, p¡0.01)—revealing participants’ intent to adjust these criteria to the task. Figure 9 shows how participants reached their final criteria in each condition. All the criteria created during the study are included in the Supplementary Materials.
On average, participants took at least one suggested improvement from 31.3% (SD=23.5%) of the automatic criteria reviews, and 78.6% (SD=20.4%) of each participant’s final criteria were revised through suggestions. Similar to how the evaluation assistant helped participants think of how to improve their prompts, the criteria review tool helped participants think of what to evaluate when they were “too busy working on the prompt” (P4). Participants mentioned how the tool “suggested [criteria] that [they] couldn’t think about” (P11) and helped them when they “didn’t really know what [they] need” (P5). For example, the review tool suggested splitting P11’s “Engagingness” criteria into “Simplicity” and “Creativity” which helped them realize the dependencies but also differences between these aspects (Tab. 2). Furthermore, participants used the review tool to refine the “questionable” (P1) criteria that they composed in order to match the quality of the criteria defined by prior research. For example, P1 used the suggestions to further refine and elaborate on their “Explainability” criteria (Tab. 2),. Through these system-supported revisions, participants felt that they had increased the overall quality of the criteria by making them more “concrete” (P1), “easy to understand” (P8), “sharper”, and “consistent” (P3). Participants noted how these improved sets “could be useful when talking with others” (P1) or even when manually evaluating outputs (P8).
@@ 7.2.3. Defining Criteria
P2 |
Merge |
Child-Friendly Language: The response should be structured in a way that promotes readability for a young child. It should use sentence structure and vocabulary appropriate for a young child’s understanding. Child-Friendly Understandability: Judge whether the response is understandable for a young child. It should use sentence structure and vocabulary appropriate for a young child’s understanding. |
@@
Simplicity: The response should use simple language and concepts that a young child can understand. Creativity: The response should include creative elements, such as analogies or stories, to make the concept interesting for a young child. |

### v2

@@ 6.2.1. Procedure
We sampled two evaluations by the Specific- Criteria condition for each of the 19 tasks, resulting in a total of 38 output pairs evaluated and 194 criteria-wise evaluations. Each evaluation was annotated for the presence of errors regarding five criteria: (1) logical (i.e., the explanation presents logical and coherent arguments and justifications), (2) faithful (i.e., the explanation does not hallucinate content that does not exist in the outputs), (3) independent (i.e., the explanation does not assess other criteria or aspects not described in the evaluation criterion), (4) evidential (i.e., the evidence extracted is relevant to the explanation), and (5) score aligned (i.e., the final score aligns with the explanation provided). We recruited two annotators, who had previous experience grading written assignments, and instructed them to mark these errors even if only part of the explanation presented the error. Since there is a large class imbalance (i.e., most explanations have no errors), we considered an explanation to have errors if at least one of the evaluators marked an error.
@@ 6.2.2. Results
Overall, the explanations were mostly free of issues: 91.4% of the explanations were logical, 99.1% were faithful, 84.2% were independent, 100% provided relevant evidence, and 98.6% were aligned with their scores. We qualitatively reviewed erroneous cases and observed that the LLM’s explanations frequently failed to be independent as they would contrast outputs on their level of detail despite the criterion not assessing this. In terms of logic, the evaluations struggled to assess creativity and could be too superficial in their interpretations. For example, news article headlines were considered to adequately address ethical dilemmas by simply including the phrase “ethical concerns”. For faithfulness, the evaluations struggled to accurately measure the length of the outputs. Finally, for score alignment, the explanations would occasionally provide one output with a higher score despite not mentioning this in its explanations. These results show that GPT-4 is able to produce relatively sensible explanations for its evaluations, but can be limited by bias towards detail and logical capabilities. While research in NLP has focused largely on improving the ”accuracy” of the scores provided by LLM-based evaluations, this suggests that more investigation is needed into the explanations provided by these evaluations.
@@ 7. User Study
To understand how the EvalLM affects the prompt iteration process when compared to following designers’ current practice, we conducted a within-subjects study where we compared EvalLM to a baseline where participants manually evaluated outputs. In this study, we aimed to answer the following research questions:
RQ1. Can EvalLM aid designers in deciding on how to revise their prompts and in verifying the effectiveness of these revisions?
@@ 7.1.1. Participants
We recruited 12 participants through posts on online forums within our institution. All participants reported having extensive experiences with prompting: nine had designed prompts for research-based applications, one designed prompts for toy projects, and two described frequent use of LLMs for productivity. Regarding the length of their experiences, four had between 1 to 3 months of experience, three had 3 to 6 months, four had 6 to 12 months, and two had 1 to 2 years. Participants were compensated with approximately 60 USD (80,000 KRW) for the 2-hour study.
@@ 7.1.2. Conditions
During the study, participants designed prompts and evaluated outputs for two given tasks. For each task, participants used EvalLM in one of two conditions: Assist or Manual. The Assist condition was the full EvalLM interface, while the Manual condition was the EvalLM interface without the evaluation assistant or the criteria review tool. In the Manual condition, participants defined their own criteria or selected from the dictionary, and then evaluated data samples by choosing which output won for each criterion. This condition supports designers’ common practices—according to our formative interviews—where they copy generated outputs into a spreadsheet to manually check which criteria were satisfied by each output.
@@ 7.1.3. Tasks
Participants designed prompts for the same two tasks. As our work focuses on novel generative tasks, we adapted tasks from two recently proposed LLM-powered HCI systems: (1) write an example that can explain a piece of scientific information to a young child—based on Lee et al.’s DAPIE system (Lee et al., 2023)—and (2) ideate a list of alternative angles for a news story—based on Peridis et al.’s AngleKindling system (Petridis et al., 2023). We chose these as they target a specific user population (i.e., children and reporters) and no significant expertise is needed to understand the task outputs. Additionally, these tasks differ in terms of their goal (i.e., explaining vs. brainstorming), and how the output relates to the input (i.e., transforming vs. expanding on the information).
@@ 7.1.4. Procedure
Participants signed the informed consent form prior to the study. After a brief introduction, participants answered a pre-task survey. After a walkthrough of the first interface, participants used this interface to design a prompt for the first task for 35 minutes.444The order of conditions and tasks were counterbalanced to mitigate ordering effects. Participants were asked to envision themselves as a developer at a startup building an application for the given task. Their goal was to design a prompt that performed better than an initial prompt designed by their team and demonstrate its performance on diverse data samples. Participants could flexibly decide on the criteria that they would be evaluating, but were asked to ensure that their final criteria set was: (1) exhaustively comprehensive (i.e., assess all factors that are important for the task), (2) mutually exclusive (i.e., minimal redundancies between criteria), and (3) clear (i.e., clearly described for others to understand what is assessed). After the task, they responded to a post-task survey and, through a semi-structured interview, we asked them about how they defined their criteria, evaluated outputs, and revised their prompts during this task. Then, participants were provided with a walkthrough of the second interface and used this interface to perform the second task for 35 minutes. After responding to the post-task survey and interview, we concluded the study with a semi-structured interview about the differences between participants’ experiences in the first and second tasks.
@@ 7.1.5. Measures
For qualitative data, we transcribed the semi-structured interviews and coded them through a thematic analysis. For quantitative data, we analyzed participants’ responses to the two post-task surveys. These surveys asked participants to rate, on a seven-point Likert scale, their self-confidence in designing prompts and evaluating outputs, and their self-perceived experience with the system based on questions from Wu et al. (Wu et al., 2022). We also asked participants to rate their own final criteria on how exhaustively comprehensive, mutual exclusive, and clear they were. Finally, participants rated their self-perceived workload on five questions from the NASA-TLX questionnaire, excluding the “Physical Demand” question. We include the detailed survey questions in Appendix D. For these Likert scale ratings, we analyzed them through the non-parametric Wilcoxon signed-rank test.
Additionally, we analyzed participants’ interaction logs to measure the number of (1) unique prompts tested, (2) criteria changes (e.g., edit, add, delete), and (3) unique outputs evaluated, where partial evaluations (i.e., only a subset of criteria were evaluated) were counted equally. For all these measures, we conducted a Shapiro-Wilk test to determine if the data was parametric, and then used a paired t-test (if parametric) and a Wilcoxon signed-rank test (if non-parametric). As participants in the study could flexibly set their own goals and requirements for their prompts, we did not conduct an external evaluation of the prompts as external evaluators may consider quality aspects that differ from participants’ intentions due to the subjectivity involved. Also, while external evaluators could be tasked to assess prompts on the criteria defined by participants, a single participant could produce criteria that differ in clarity and complexity across tasks and conditions, which could lead to unfair comparisons.
7.2. Results
In this section, we describe findings on how participants evaluated outputs and revised their prompts in §7.2.1 and §7.2.2 (RQ1), how they defined their criteria in §7.2.3 (RQ2), their trust for the evaluation assistant in §7.2.4 (RQ3), and their overall perceived workload in §7.2.5. For each of these, we first describe relevant quantitative findings and then qualitative insights.
7.2.1. Evaluating Outputs
Overall, participants had higher self-confidence in their ability to evaluate prompt outputs with the Assist condition (Assist=6.71 ± 0.40, Manual=4.96 ± 0.72, z=9.23, p¡0.001). Participants explained that the Assist condition shouldered the burden of examining outputs, which enabled them to evaluate prompts more comprehensively. Specifically, with Assist, participants were able to evaluate a larger number of unique outputs (Assist=20.42 ± 14.46, Manual=10.08 ± 6.27, z=8.00, p=0.03). Several participants noted how this helped them look at the “bigger picture” (P1) and understand “how [a prompt] will work at a larger scale” (P5).
By facilitating evaluations, the Assist condition also allowed participants to evaluate more criteria and assess performance on diverse dimensions, without worrying about the cost involved. For example, P9 mentioned that they were encouraged to “think of more aspects that they wanted to evaluate” and P2 mentioned that they “just kept criteria […] to check them just in case.” In contrast, when using the Manual condition, participants would frequently evaluate samples only on a subset of their criteria—on average, 46.7% (SD=31.2%) of all evaluated samples were partially evaluated. Besides automating evaluation, several participants also mentioned how the Assist condition made it easier to evaluate the outputs themselves. While it could be challenging “tell what is different by simply looking at [outputs]” (P5), the evaluation assistant’s explanations and highlighting in outputs made it “easier to tell outputs apart” (P6).
As the Assist condition supported faster evaluation cycles, participants mentioned how they used the evaluation assistant as a “debugger” (P8). The assistant helped participants to easily check “what the prompt is already satisfying” (P8) and “where it’s lacking” (P9). P11 mentioned, “The [system] tells me that my prompt is worse on simplicity so that means that my prompt is not communicating this clearly and I should focus on fixing this.” With the evaluation support, participants could quickly identify aspects or “bugs” that needed to be handled and, after revising their prompts, quickly verify whether these issues had been handled or not.
@@ 7.2.2. Revising Prompts
Besides aiding participants in checking revisions, the Assist condition also helped participants in thinking about how to revise their prompts. Participants felt that the Assist condition helped them think about how to complete the task better (Assist=6.83 ± 0.39, Manual=5.67 ± 1.15, z=0.00, p=0.01) and that they were collaborating with the system to design the prompts (Assist=6.58 ± 0.51, Manual=4.58 ± 1.98, z=1.50, p¡0.01) (Fig. 8, left). While participants also reported a higher self-confidence on their ability to design and improve prompts with Assist, this difference was not significant (Assist=5.63 ± 1.03, Manual=5.00 ± 0.95, z=11.00, p=0.17). Additionally, we found that participants with the Assist condition reached a similar level of satisfaction in their prompts (Assist=5.67 ± 1.61, Manual=5.17 ± 1.47, z=20.00, p=0.24) (“Match Goal” in Fig. 8) by testing a smaller number of prompt variations (Assist=5.00 ± 3.13, Manual=8.42 ± 4.76, z=1.50, p=0.01).
@@
| P2 | Merge |
Child-Friendly Language: The response should be structured in a way that promotes readability for a young child. It should use sentence structure and vocabulary appropriate for a young child’s understanding.
@@
Simplicity: The response should use simple language and concepts that a young child can understand.
@@
In the Manual condition, participants mentioned how they felt like “the only brain thinking about [the task]” (P10). In contrast, participants felt that Assist condition was providing them with “feedback on what to improve” (P2) and could help them set the “direction” for their prompts (P7, P9). Several participants found the explanations to be particularly useful as they presented them with “diverse opinions” (P1) that they “couldn’t think about” (P6). For example, P5 purposefully increased the number of evaluation trials and checked the explanation for each trial to “see all the differences [in opinion] and then change [their] prompt to satisfy all of these.” In this sense, the Assist condition could help participants gain a more holistic understanding of how to improve their prompts and lead them to more effective prompt revisions—as reflected by how fewer prompt changes were tested in the Assist condition.
These explanations, however, could also lead participants to feel less in control and more aware of their prompt’s weaknesses. P2, the only participant that preferred the Manual condition, explained how she could incorporate her “own diverse ideas” into her prompt in Manual but, in Assist, she “kept worrying about the evaluations” and would predominantly focus on incorporating the LLM’s feedback. Even participants that preferred the Assist condition noted limitations in how the evaluations frequently returned ties or high ratings to both outputs, without indicating what could be improved. Since the feedback focused on the outputs rather than the prompts, several participants expressed how they wanted the LLM to also “automatically suggest ways to improve prompts” (P3).
7.2.3. Defining Criteria
Participants finished the tasks with a similar number of criteria in both conditions (Assist=4.25 ± 1.29, Manual=4.00 ± 1.28, t=10.50, p=0.55). However, participants felt that their criteria in the Assist condition were more comprehensive (Assist=6.00 ± 1.48, Manual=4.75 ± 1.48, t=2.07, p=0.06), although with marginal significance, and clearer (Assist=6.42 ± 0.67, Manual=4.92 ± 1.44, t=3.59, p¡0.01) (Fig. 8, right). Additionally, participants made more changes to their criteria with the Assist condition (Assist=22.67 ± 9.17, Manual=13.33 ± 8.64, t=2.29, p=0.04). Regarding participants’ final criteria, there was no significant difference in the proportion of criteria created from scratch (Assist=30.1% ± 34.5%, Manual=16.67% ± 18.8%, w=15.0, p=0.37) or through the criteria dictionary (Assist=77.78% ± 20.5%, Manual=69.93% ± 34.50%, w=22.0, p=0.95). However, in Assist, a significantly higher proportion of the dictionary-based criteria had been edited (Assist=75.8% ± 24.1%, Manual=17.4% ± 34.5%, w=4.000, p¡0.01)—revealing participants’ intent to adapt these criteria to specific tasks. Figure 9 shows how participants reached their final criteria.
On average, participants took at least one suggested improvement from 31.3% (SD=23.5%) of the automatic criteria reviews, and 78.6% (SD=20.4%) of their final criteria were revised through suggestions. Similar to how the evaluation assistant helped participants think of how to improve prompts, the criteria review tool helped them think of what to evaluate by “suggesting [criteria] that [they] couldn’t think about” (P11) and helping them when they “didn’t really know what [they] need” (P5). For example, P11 received a suggestion to split “Engagingness” into “Simplicity” and “Creativity” which helped them realize the dependency and difference between these aspects (Tab. 2). Furthermore, participants used the review tool to refine any “questionable” (P1) criteria that they wrote themselves in order to match the quality of criteria definitions from prior research. Through these system-supported revisions, participants increased the overall quality of their criteria where they could be useful when “communicating with others” (P1) or even when manually evaluating outputs (P8).